OpenStudy (anonymous):

The absolute value of 2/3x = 8 How do I solve it?

geerky42 (geerky42):

|2/3x| = 8 2/3x = ±8

OpenStudy (anonymous):

A. {12, 24} B. {–12, –24} C. {12, –12} D. {–12, –12

OpenStudy (anonymous):

\[\frac{2}{3}x=8\iff x=8\times \frac{3}{2}=4\times 3=12\] for the first one

OpenStudy (anonymous):

\[\frac{2}{3}x=-8\iff x=-8\times \frac{3}{2}=-12\] for the second

OpenStudy (anonymous):

Thank you :D The teacher didn't explain it so well